OpenStudy (nannymcphee):

whats the vertex of -x^2-7x-12

OpenStudy (calculusxy):

vertex of the parabola is either the max. or min. value of the parabola. you need to first find the axis of symmetry: x = -x/2a after finding that x value, plug it back into the equation -x^2-7x-12.

OpenStudy (calculusxy):

sorry correction: axis of symmetry = \(\large \frac{-b}{2a}\)

OpenStudy (anonymous):

Hi again ^_^ Do you know the formula of finding the vertex of a quadratic equation? Let's start with the general Quadratic Equation: Ax^2 + Bx +C=0 Now let's look at your Equation: -x^2-7x-12 So this means that your A= -1, B = -7 and C= -12 Follow what calculuxy posted, that is the formula for the vertex :)

OpenStudy (calculusxy):

therefore evaluate: \(\large \frac{-(-7)}{2(-1)}\) and you should get the axis of symmetry